{2w+5y=-24
{3w-5y=14 please help me with this problem

Answer by elima(1433)

{2w+5y=-24
{3w-5y=14
Add both equations together to eliminate one variable;

2w+5y=-24
3w-5y=14
----------
5w=-10
Divide both sides by 5;
w=-2
Now that you know what w = we can plug that in the equation and find y;
2(-2)+5y=-24
-4+5y=-24
5y=-24+4
5y=-20
y=-4
w=-2
y=-4
Hope you understand
=)